Currie Company borrowed $29,000 from the Sierra Bank by issuing an 11% three-year note. Currie agreed to repay the principal and interest by making annual payments in the amount of $11,867. Based on this information, the amount of the interest expense associated with the second payment would be: (Round your answer to the nearest
Answer #1
Date Cash payment Interest expense Principal payment Ending balance of Note
1st year beginning 29,000
1st year end 11,867 29,000 x 11% = 3,190 11867-3,190 = 8,677 29,000-8,677 = 20,323
2nd year end 11,867 20,323 x 11% = 2,236 11,867-2,236 = 9,631 20,323-9,631 = 10,692

Interest expense associated with second payment = $2,236

If interest expense is rounded to nearest cent, then interest expense associated with second payment will be $2,235.53.

Kindly comment if you need further assistance. Thanks‼!